# Kai Yao    Marc Juarez

### v1

@@ VI Results
This section presents the results of our security evaluation. We organize it into five parts: first, we assess MFD technique robustness to removal across threat models and attack strategies; second, we repeat the evaluation for forgery; third, we study the relationship between removal and forgery; fourth, we test whether higher utility of the MFD technique is an indicator of vulnerability to the attacks; and, fifth, we investigate the impact of model’s type and architecture on ASR.
@@ VI-A Robustness to Removal
TABLE II: Attack success rates (ASR, mean $\pm$ std, in %) of removal attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1) and fingerprint-agnostic perturbations (B2: Gaussian noise, blur, JPEG compression, resizing). Higher ASR indicates greater vulnerability. The maximum values per method are highlighted in bold. The row separators indicate the domains in Table I: RGB, Frequency, and Learned.
@@
McCloskey18 [McCloskey18] 36.70 $\pm$ 5.16 59.36 $\pm$ 28.81 13.49 $\pm$ 0.55 2.09 $\pm$ 1.00 13.33 $\pm$ 5.70 2.88 $\pm$ 1.75 11.97 $\pm$ 5.89
Nataraj19 [Nataraj19] 97.53 $\pm$ 1.84 86.54 $\pm$ 12.49 84.75 $\pm$ 0.94 27.97 $\pm$ 0.85 38.31 $\pm$ 3.16 46.16 $\pm$ 4.09 51.42 $\pm$ 1.14
Nowroozi22 [Nowroozi22] 84.94 $\pm$ 1.32 67.87 $\pm$ 1.75 84.55 $\pm$ 0.97 22.47 $\pm$ 1.69 29.00 $\pm$ 1.26 32.43 $\pm$ 1.66 42.39 $\pm$ 2.87
Song24-RGB [Song24] 91.38 $\pm$ 2.63 2.34 $\pm$ 0.61 0.00 $\pm$ 0.00 1.19 $\pm$ 0.40 0.18 $\pm$ 0.32 0.00 $\pm$ 0.00
Marra19a [Marra19a] 100.00 $\pm$ 0.00 40.71 $\pm$ 5.20 9.55 $\pm$ 2.10 1.67 $\pm$ 1.07 13.50 $\pm$ 9.26 4.10 $\pm$ 1.06 5.04 $\pm$ 1.18
Durall20 [Durall20] 99.90 $\pm$ 0.20 95.23 $\pm$ 7.93 99.46 $\pm$ 0.94 20.83 $\pm$ 5.80 61.63 $\pm$ 2.21 39.04 $\pm$ 4.51 94.12 $\pm$ 1.11
Dzanic20 [Dzanic20] 98.90 $\pm$ 1.71 99.10 $\pm$ 1.37 99.78 $\pm$ 0.22 40.33 $\pm$ 0.18 90.59 $\pm$ 4.08 48.26 $\pm$ 0.54 98.63 $\pm$ 0.61
Giudice21 [Giudice21] 100.00 $\pm$ 0.00 92.74 $\pm$ 8.01 21.99 $\pm$ 1.30 59.21 $\pm$ 3.61 45.48 $\pm$ 1.80 81.18 $\pm$ 1.76
Corvi23-R [Corvi23] 100.00 $\pm$ 0.00 35.16 $\pm$ 3.84 2.44 $\pm$ 1.03 17.02 $\pm$ 4.26 1.11 $\pm$ 0.37 25.20 $\pm$ 6.53
Corvi23-S [Corvi23] 77.44 $\pm$ 27.47 15.08 $\pm$ 7.42 0.42 $\pm$ 0.72 27.90 $\pm$ 18.07 0.42 $\pm$ 0.72 32.21 $\pm$ 15.68
Song24-Freq [Song24] 31.08 $\pm$ 3.82 70.56 $\pm$ 6.94 5.30 $\pm$ 0.87 50.55 $\pm$ 3.78 6.79 $\pm$ 1.55 68.64 $\pm$ 0.63
Qian20 [Qian20] 90.71 $\pm$ 6.43 82.05 $\pm$ 6.12 56.08 $\pm$ 38.06 67.12 $\pm$ 28.49 26.28 $\pm$ 6.65 75.90 $\pm$ 20.89
Wang20 [Wang20] 100.00 $\pm$ 0.00 78.19 $\pm$ 18.90 0.67 $\pm$ 0.77 37.08 $\pm$ 54.53 34.17 $\pm$ 57.01 49.04 $\pm$ 44.23
Song24-SL [Song24] 84.45 $\pm$ 0.85 55.86 $\pm$ 1.97 2.68 $\pm$ 0.74 20.18 $\pm$ 1.66 7.77 $\pm$ 1.51 34.61 $\pm$ 2.59
@@
Removal attacks are broadly effective across most MFD methods. In many cases, the strongest attack achieves over $80\%$ ASR under the $\ell_{\infty}$ norm constraint, indicating that fingerprints can often be suppressed while maintaining high perceptual quality. Corresponding LPIPS and PSNR values remain within good quality ranges for most attacks, and representative attacked images confirm that the perturbations are imperceptible (see Tables IV, V, and Figure 5 in Appendix A).
The main exceptions are Marra19a and Song24-RGB, which exhibit strong resistance to most attacks unless white-box access is available. McCloskey18, Corvi23-R, and Corvi23-S also show some resistance to black-box attacks; however, their baseline (non-adversarial) accuracies are substantially lower. Marra19a relies on noise residuals (image minus its BM3D-denoised version), and our results indicate that structured residuals are difficult to fully eliminate under black-box settings, making the method more robust. Song24-RGB measures the distance of a sample to the true data manifold in the RGB domain. Because fingerprints from the same model likely form tight clusters in this space and our attacks enforce an $\ell_{\infty}$ perturbation bound in the RGB domain, perturbed samples rarely escape their cluster, limiting the effectiveness of removal. In contrast, Song24-Freq and Song24-SL lack such constraints in the frequency or learned feature domains, which explains their greater vulnerability to black-box attacks compared to Song24-RGB. Since these methods are highly vulnerable to white-box attacks yet robust in black-box scenarios, it is likely that their fingerprints are inherently less transferable from the image space. Taken together, these findings suggest that noise residual- and manifold-based representations holds promise for robustness against removal in black-box scenarios.
Table II shows that ASR varies based on adversarial knowledge. White-box attacks (W1–W3) generally outperform black-box attacks (B1 and B2) across MFD techniques, which is expected since white-box attackers have privileged visibility into the MFD system. Among white-box, W1 attacks consistently achieve the best performance—when they are applicable. However, W1 attacks are not applicable for techniques that do not support direct gradient optimization. For these cases, W2 and W3 attacks rely on approximations of the gradients. As seen in the table, these approximations sometimes underperform relative to B1 strategies, which are also approximated. This counterintuitive result arises when the approximation errors of W2 and W3 exceed those of B1.
@@
Although adaptive strategies (W1–W3, and B1) always outperform fingerprint-agnostic strategies based on simple image manipulations (B2), these generic black-box perturbations alone often achieve ASRs $50\%$ and achieve ASR $90\%$ for one of the techniques (Dzanic20), indicating that fingerprint traces can be significantly disrupted without any knowledge of the underlying MFD technique. This further demonstrates the fragility of current MFD methods and heightens concerns about the applicability of these techniques in adversarial settings where not even costly attacks are needed.
ASRs show low standard deviation across most attack instances. However, some attacks exhibit high variance, with W3 against McCloskey18 showing up to $28.81$ pp. difference in ASR. The MFD technique appears to be the primary driver of this variance, creating uncertainty for users about attack robustness. This variability suggests certain samples are inherently more vulnerable to fingerprint removal. We leave a detailed investigation of the underlying causes for future work.
@@
McCloskey18 [McCloskey18] 7.51 $\pm$ 2.82 10.43 $\pm$ 2.81 2.10 $\pm$ 0.06
Nataraj19 [Nataraj19] 51.64 $\pm$ 2.78 9.85 $\pm$ 3.17 11.27 $\pm$ 0.72
Nowroozi22 [Nowroozi22] 12.09 $\pm$ 3.57 6.84 $\pm$ 0.73 10.20 $\pm$ 1.57
Song24-RGB [Song24] 16.42 $\pm$ 4.59 0.00 $\pm$ 0.00
Marra19a [Marra19a] 97.85 $\pm$ 2.90 4.46 $\pm$ 1.28 0.50 $\pm$ 0.52
Durall20 [Durall20] 57.67 $\pm$ 8.63 23.49 $\pm$ 2.70 18.18 $\pm$ 1.13
Dzanic20 [Dzanic20] 15.19 $\pm$ 5.09 25.53 $\pm$ 9.61 25.51 $\pm$ 3.93
Giudice21 [Giudice21] 98.92 $\pm$ 0.39 13.95 $\pm$ 4.09
Corvi23-R [Corvi23] 99.77 $\pm$ 0.40 3.67 $\pm$ 1.38
Corvi23-S [Corvi23] 9.12 $\pm$ 4.06 2.52 $\pm$ 1.47
Song24-Freq [Song24] 3.31 $\pm$ 1.38 6.66 $\pm$ 0.88
Qian20 [Qian20] 18.12 $\pm$ 15.79 12.51 $\pm$ 1.64
Wang20 [Wang20] 99.55 $\pm$ 0.77 5.68 $\pm$ 5.09
Song24-SL [Song24] 10.09 $\pm$ 1.90 7.29 $\pm$ 1.47
@@ VI-B Robustness to Forgery
Table III shows that forgery ASRs are below those of removal across the board. Forgery is a fundamentally harder task than removal: with 12 models in the attribution pool, successful forgery requires targeted misattribution to a specific model, while removal only requires misclassification into any incorrect class. Geometrically, removal corresponds to crossing the nearest decision boundary, whereas forgery demands reaching a potentially narrow and distant target region. When only two models are considered (which is the case for general deepfake detection tasks), the two objectives coincide and removal corresponds to forgery, but as the number of models grows, the decision landscape becomes increasingly complex and harder to manipulate by a forgery attack.
@@
Despite this difficulty, many methods still exhibit substantial forgery ASRs, especially in white-box settings. On average, W1 yields the highest ASRs, followed by W2 and then W3, consistent with their decreasing levels of gradient access—from direct gradients (W1) to analytic approximations (W2) to surrogate-based approximations (W3). Even in the black-box setting, ASRs reach up to 25.51%, indicating that a non-trivial fraction of samples can still be intentionally misattributed to a chosen target model. Importantly, note that the reported means and standard deviations are computed by averaging over all 12 generative models. Refer to Figure 4 for the variation in attack vulnerability of each model when used as a forgery target. In practice, the attacker may be interested in a specific model, or a specific model family that might be more vulnerable to attacks, in which case the average ASR within the family might be significantly higher. Furthermore, since our gradient-based attacks are generic and not tailored to the internal design of each method, method-specific adaptations could plausibly yield even higher forgery success rates.
@@
Similar to removal attacks, forged samples cannot be visually distinguished from their original counterparts, maintaining an LPIPS <0.05<0.05 and PSNR >35>35 dB across all forged samples (see Tables VI, VII in Appendix A). It is important to note that higher ASR for forgery and removal could be achieved if the adversary relaxed our strict $\ell_{\infty}$ bound constraints, as explained in Appendix D.
@@
Recall that we discarded B2 attack strategies from the forgery attack evaluation because they lack specificity to embed a targeted model fingerprint into the forged sample. For the remaining attacks, we observe trends similar to those in the removal evaluation. As with removal, white-box settings tend to outperform B1 strategies, with W1 consistently achieving the highest ASRs. In contrast, ASR variance for McCloskey18 is relatively lower than in removal attacks, while Qian20 now exhibits the highest variance. This indicates that while the objectives of removal and forgery are similar, they are not perfectly complementary—a finding that motivates a deeper investigation of their relationship.
@@
We observe a clear positive correlation between the ASR of removal and forgery, for both white-box and black-box scenarios. As shown in Fig. 2, MFD methods that are more susceptible to removal are also generally more vulnerable to forgery. This trend is intuitive: since all reported attacks are gradient-based, if gradients can efficiently guide perturbations toward a distant forgery region, they can likewise push samples outside their current decision region, enabling removal.
@@
Figure 3: Trade-off between attribution utility (accuracy of the classifier $h$, y-axis) and robustness (ASR, x-axis) across four scenarios: removal under white-box (far left) and black-box I (second), and forgery under white-box (third) and black-box I (far right). White-box scenario uses the best ASR per method (highest ASR among W1, W2, and W3). Each point is an MFD method; error bars show standard deviation across runs. Methods nearer the top-left deliver stronger utility–robustness trade-offs. Green circles indicate outliers with better trade-off. See Figure 7 in Appendix C for individual whit-box attack data.
However, for the black-box scenario, this correlation does not imply that high removal ASR translates into high forgery ASR in absolute value terms. For B1 (Figure 2, Right), we observed that the increase in forgery ASR from less-resilient to more-resilient methods remains small in magnitude. The magnitude of this increase likely depends on the size of the model pool and the characteristics of the models themselves. For instance, in the simplest case of a two-model pool, the attribution problem reduces to binary classification (similar to deepfake detection). In this setting, removing fingerprints of one model effectively forces the sample to be attributed to the other, making removal and forgery effectively the same goal. By contrast, as the number of models in the pool grows, removal remains comparably feasible—since pushing samples outside the decision region of the true model does not fundamentally depend on the pool size and other models—while forgery can become increasingly difficult. This is because forging requires moving a sample into a specific target region, and the difficulty of reaching such regions grows with the number and arrangement of models’ decision boundaries.
Hence, both removal and forgery ASR scales can be affected by the geometric landscape of decision regions across models in each individual model attribution task. The precise relationship between removal and forgery thus depends on the size of the model pool and the characteristics of the models. A more detailed analysis of how pool size and model diversity govern this relationship is left to closer examination.

### v2

@@ VI Results
This section presents the results of our security evaluation. We organize it into five parts: first, we assess MFD technique robustness to removal across threat models and attack strategies; second, we repeat the same evaluation for forgery; third, we study the relationship between removal and forgery; fourth, we test whether higher utility of the MFD technique is an indicator of attack success; and, fifth, we investigate the impact of the model’s type and architecture on ASR.
@@ VI-A Robustness to Removal
TABLE II: Attack success rates (ASR, mean $\pm$ std, in %) of removal attacks against evaluated MFD methods. White-box attacks include direct gradient-based optimization (W1), analytic approximations (W2), and surrogate fingerprint extractors (W3). Black-box attacks include surrogate classifiers (B1) and fingerprint-agnostic perturbations (B2: Gaussian noise, blur, JPEG compression, resizing). Higher ASR indicates greater vulnerability. The maximum values per method are highlighted in bold. The row separators indicate the domains: RGB, Frequency, and Learned.
@@
McCloskey18 [McCloskey18] $36.70$ p m 5.16 $59.36$ p m 28.81 $13.49$ p m 0.55 $2.09$ p m 1.00 $13.33$ p m 5.70 $2.88$ p m 1.75 $11.97$ p m 5.89
Nataraj19 [Nataraj19] $97.53$ p m 1.84 $86.54$ p m 12.49 $84.75$ p m 0.94 $27.97$ p m 0.85 $38.31$ p m 3.16 $46.16$ p m 4.09 $51.42$ p m 1.14
Nowroozi22 [Nowroozi22] $84.94$ p m 1.32 $67.87$ p m 1.75 $84.55$ p m 0.97 $22.47$ p m 1.69 $29.00$ p m 1.26 $32.43$ p m 1.66 $42.39$ p m 2.87
Song24-RGB [Song24] $91.38$ p m 2.63 $2.34$ p m 0.61 $0.00$ p m 0.00 $1.19$ p m 0.40 $0.18$ p m 0.32 $0.00$ p m 0.00
Marra19a [Marra19a] $100.00$ p m 0.00 $40.71$ p m 5.20 $9.55$ p m 2.10 $1.67$ p m 1.07 $13.50$ p m 9.26 $4.10$ p m 1.06 $5.04$ p m 1.18
Durall20 [Durall20] $99.90$ p m 0.20 $95.23$ p m 7.93 $99.46$ p m 0.94 $20.83$ p m 5.80 $61.63$ p m 2.21 $39.04$ p m 4.51 $94.12$ p m 1.11
Dzanic20 [Dzanic20] $98.90$ p m 1.71 $99.10$ p m 1.37 $99.78$ p m 0.22 $40.33$ p m 0.18 $90.59$ p m 4.08 $48.26$ p m 0.54 $98.63$ p m 0.61
Giudice21 [Giudice21] $100.00$ p m 0.00 $92.74$ p m 8.01 $21.99$ p m 1.30 $59.21$ p m 3.61 $45.48$ p m 1.80 $81.18$ p m 1.76
Corvi23-R [Corvi23] $100.00$ p m 0.00 $35.16$ p m 3.84 $2.44$ p m 1.03 $17.02$ p m 4.26 $1.11$ p m 0.37 $25.20$ p m 6.53
Corvi23-S [Corvi23] $77.44$ p m 27.47 $15.08$ p m 7.42 $0.42$ p m 0.72 $27.90$ p m 18.07 $0.42$ p m 0.72 $32.21$ p m 15.68
Song24-Freq [Song24] $31.08$ p m 3.82 $70.56$ p m 6.94 $5.30$ p m 0.87 $50.55$ p m 3.78 $6.79$ p m 1.55 $68.64$ p m 0.63
Qian20 [Qian20] $90.71$ p m 6.43 $82.05$ p m 6.12 $56.08$ p m 38.06 $67.12$ p m 28.49 $26.28$ p m 6.65 $75.90$ p m 20.89
Wang20 [Wang20] $100.00$ p m 0.00 $78.19$ p m 18.90 $0.67$ p m 0.77 $37.08$ p m 54.53 $34.17$ p m 57.01 $49.04$ p m 44.23
Song24-SL [Song24] $84.45$ p m 0.85 $55.86$ p m 1.97 $2.68$ p m 0.74 $20.18$ p m 1.66 $7.77$ p m 1.51 $34.61$ p m 2.59
@@
Removal attacks are broadly effective across most MFD methods. In many cases, the strongest attack achieves over $80\%$ ASR under the $\ell_{\infty}$ norm constraint, indicating that fingerprints can often be suppressed while maintaining high perceptual quality. Corresponding LPIPS and PSNR values remain within good image quality ranges for most attacks, and upon visual inspection of representative attacked images we confirm that the perturbations are imperceptible (see Tables IV, V, and Figure 5 in Appendix A).
The exceptions are Marra19a and Song24-RGB, which exhibit strong resistance to most attacks unless white-box access is available. McCloskey18, Corvi23-R, and Corvi23-S also show some resistance to black-box attacks; however, their baseline (non-adversarial) accuracies are substantially lower. Marra19a relies on noise residuals (image minus its BM3D-denoised version), and our results indicate that structured residuals are difficult to fully eliminate under black-box settings, making the method more robust. Song24-RGB measures the distance of a sample to the true data manifold in the RGB domain. Because fingerprints from the same model likely form tight clusters in this space and our attacks enforce an $\ell_{\infty}$ perturbation bound in the RGB domain, perturbed samples rarely escape their cluster, limiting the effectiveness of removal. In contrast, Song24-Freq and Song24-SL lack such constraints in the frequency or learned feature domains, which explains their greater vulnerability to black-box attacks compared to Song24-RGB. Since these methods are highly vulnerable to white-box attacks yet robust in black-box scenarios, it is likely that their fingerprints are inherently less transferable from the image space. Taken together, these findings suggest that noise residual- and manifold-based representations holds promise for robustness against removal in black-box scenarios.
Table II shows that ASR varies based on adversarial knowledge. White-box attacks (W1–W3) generally outperform black-box attacks (B1 and B2) across MFD techniques, which is expected since white-box attackers have privileged visibility into the MFD system. Among white-box attacks, W1 consistently achieves the best performance—when applicable. However, W1 attacks are not applicable for techniques that do not support direct gradient optimization. For these cases, W2 and W3 attacks rely on approximations of the gradients. As seen in the table, these approximations sometimes underperform relative to B1 strategies, which are also approximated. This counterintuitive result arises when the approximation errors of W2 and W3 exceed those of B1.
@@
Although adaptive strategies (W1–W3, and B1) always outperform fingerprint-agnostic strategies based on simple image manipulations (B2), these generic black-box perturbations alone often achieve ASRs $50\%$, with one of the techniques (Dzanic20) being particularly vulnerable (ASR $90\%$). This result indicates that fingerprint traces can be significantly disrupted without any knowledge of the underlying MFD technique, demonstrating the fragility of current MFD methods and raising concerns about the applicability of these techniques in adversarial scenarios.
ASRs show low standard deviation across most attack instances. However, some attacks exhibit high variance, with W3 against McCloskey18 showing up to $28.81$ pp. difference in ASR. The choice of MFD technique appears to be the main driver of this variance, creating uncertainty for users about attack robustness. This variability suggests certain samples are inherently more vulnerable to fingerprint removal. We leave a detailed investigation of the underlying causes for future work.
@@
McCloskey18 [McCloskey18] $7.51$ p m 2.82 $10.43$ p m 2.81 $2.10$ p m 0.06
Nataraj19 [Nataraj19] $51.64$ p m 2.78 $9.85$ p m 3.17 $11.27$ p m 0.72
Nowroozi22 [Nowroozi22] $12.09$ p m 3.57 $6.84$ p m 0.73 $10.20$ p m 1.57
Song24-RGB [Song24] $16.42$ p m 4.59 $0.00$ p m 0.00
Marra19a [Marra19a] $97.85$ p m 2.90 $4.46$ p m 1.28 $0.50$ p m 0.52
Durall20 [Durall20] $57.67$ p m 8.63 $23.49$ p m 2.70 $18.18$ p m 1.13
Dzanic20 [Dzanic20] $15.19$ p m 5.09 $25.53$ p m 9.61 $25.51$ p m 3.93
Giudice21 [Giudice21] $98.92$ p m 0.39 $13.95$ p m 4.09
Corvi23-R [Corvi23] $99.77$ p m 0.40 $3.67$ p m 1.38
Corvi23-S [Corvi23] $9.12$ p m 4.06 $2.52$ p m 1.47
Song24-Freq [Song24] $3.31$ p m 1.38 $6.66$ p m 0.88
Qian20 [Qian20] $18.12$ p m 15.79 $12.51$ p m 1.64
Wang20 [Wang20] $99.55$ p m 0.77 $5.68$ p m 5.09
Song24-SL [Song24] $10.09$ p m 1.90 $7.29$ p m 1.47
@@ VI-B Robustness to Forgery
Table III shows that forgery ASRs are below those of removal across the board. Forgery is a fundamentally harder task than removal: with 12 models in the pool, successful forgery requires targeted misattribution to a specific model, while removal succeeds upon misclassification into any incorrect class. Geometrically, removal corresponds to crossing the nearest decision boundary, whereas forgery requires pushing the image to a potentially farther target region. When MFD considers only two classes, such as in deepfake detection, removal and forgery become equivalent: erasing traces of one source implies attribution to the other. But as the number of classes grows, the decision landscape becomes increasingly complex, making forgery attacks harder to execute.
@@
Despite this difficulty, many methods still exhibit substantial forgery ASRs, especially in white-box settings. On average, W1 yields the highest ASRs, followed by W2 and then W3, consistent with their decreasing levels of gradient access—from direct gradients (W1) to analytic approximations (W2) to surrogate-based approximations (W3). Even in the black-box setting, ASRs reach up to 25.51%, indicating that a non-trivial fraction of samples can still be intentionally misattributed to a chosen target model. Importantly, note that the reported means and standard deviations are computed by averaging over all 12 generative models. Refer to Figure 4 for the variation in attack vulnerability of each model when used as a forgery target. In practice, the attacker may be interested in a specific model, or a specific model family that might be more vulnerable to attacks, in which case the average ASR within the family might be significantly higher. Furthermore, our gradient-based attacks are method-agnostic rather than exploiting the specific architecture of each detector; tailored attacks would likely achieve even higher forgery ASRs.
@@
Similar to removal attacks, forged samples cannot be visually distinguished from their original counterparts, maintaining an LPIPS <0.05<0.05 and PSNR >35>35 dB across all forged samples (see Tables VI, VII in Appendix A). It is important to note that even higher ASRs for both forgery and removal are feasible if the adversary relaxes the constraint on image quality ($\ell_{\infty}$ bound), as explained in Appendix D.
@@
Recall that we discarded B2 attack strategies from the forgery attack evaluation because they lack specificity to embed a targeted model fingerprint into the forged sample. For the remaining attacks, we observe trends similar to those in the removal evaluation. As with removal, white-box settings tend to outperform B1 strategies, with W1 consistently achieving the highest ASRs. In contrast, ASR variance for McCloskey18 is relatively lower than in removal attacks, while Qian20 now exhibits the highest variance. This finding indicates that while the objectives of removal and forgery are similar, they are not perfectly complementary, and motivates a deeper investigation of the relationship between removal and forgery.
@@
We observe a clear positive correlation between the ASR of removal and forgery, for both white-box and black-box scenarios. As shown in Fig. 2, MFD methods that are more susceptible to removal are also generally more vulnerable to forgery. This trend is intuitive: since all reported attacks are gradient-based, if gradients can efficiently guide perturbations toward a distant forgery region, they can likewise push samples outside their current decision region, leading to removal.
@@
Figure 3: Trade-off between attribution utility (accuracy of the classifier $h$, y-axis) and robustness (ASR, x-axis) across four scenarios: removal under white-box (far left) and black-box I (second), and forgery under white-box (third) and black-box I (far right). White-box scenario uses the best ASR per method (highest ASR among W1, W2, and W3). Each point is an MFD method; error bars show standard deviation across runs. Methods nearer the top-left deliver stronger utility–robustness trade-offs. Green circles indicate outliers with better trade-off. See Figure 7 in Appendix C for a more detailed analysis on white-box attacks.
However, for the black-box scenario, this correlation does not imply that high removal ASR translates into high forgery ASR in absolute value terms. For B1 (Figure 2, Right), we observed that the increase in forgery ASR from less-resilient to more-resilient methods remains small in magnitude. The magnitude of this increase likely depends on the size of the model pool and the characteristics of the models themselves. For instance, in the simplest case of a two-model pool, the attribution problem reduces to binary classification (similar to deepfake detection). In this setting, removing fingerprints of one model effectively forces the sample to be attributed to the other, making removal and forgery effectively the same goal. As the model pool grows, removal remains comparably feasible because pushing samples outside the true model’s decision region does not depend on the other models. Forgery, by contrast, becomes increasingly difficult: it requires moving a sample into a specific target region, and reaching that region becomes harder with more decision boundaries.
Both removal and forgery ASR scales can be affected by the geometric landscape of decision regions across models in each individual model attribution task. The precise relationship between removal and forgery thus depends on the size of the model pool and the characteristics of the models. A more detailed analysis of how pool size and model diversity govern this relationship is left to closer examination.